# History-Aware Neural Operator: Robust Data-Driven Constitutive Modeling of Path-Dependent Materials

### v1

@@ 4.2.1 Results and discussion
For Testset II, where an initial portion of the loading path is missing, RNN1 exhibits a dramatic increase in prediction error to NRMSE=$35.9\%$, reflecting its significant reliance on early strain increments for proper hidden state initialization. RNN2, which incorporates both strain–stress history as inputs, attains improved accuracy than RNN1; however, it still experiences a substantial error increase—fivefold to $10.6\%$. In contrast, HANO maintains exceptionally high accuracy (NRMSE of $0.4\%$), nearly two orders of magnitude lower than the error of RNN1, demonstrating the superior performance of neural operator architectures over RNN-based models. In addition, the comparison between RNN1 and RNN2 highlights that using strain–stress history, rather than a single strain sequence, significantly enhances model predictivity.
@@
By incorporating both strain and stress information and employing the designed neural operator framework, the proposed HANO approach effectively overcomes the zero-stress initialization constraints that limit conventional RNN-based models. This suggests its superior capacity for learning the hidden dynamics that govern path-dependent, irreversible behavior, enabling accurate stress predictions from any intermediate point along the loading history, even when initial data segments are missing.
@@ 4.3.2 Effect of discretization resolution in testset
In contrast, the proposed HANO model maintains consistently low error across all resolution levels—$1.1\%$ at $50$ steps, $1.2\%$ at $60$ steps, $0.6\%$ at $100$ steps, and $0.7\%$ at $150$ steps—with only minimal degradation at the lowest resolutions. This highlights the resolution-invariant property of the HANO framework. Even at the lowest tested resolution ($60$ time steps), HANO achieves better performance than the RNN model at its optimal resolution ($100$ steps), with errors of $1.1\%$ versus $2\%$, demonstrating robustness to discretization variation that derives from its neural-operator backbone. The slight increase in HANO error at very coarse resolutions is attributed to path discontinuities that arise when the loading history is too sparsely sampled, introducing higher abrupt transitions that can challenge even resolution-invariant surrogates.
Figure 4: Comparison of prediction errors (NRMSE) for RNN2 and HANO approaches with varying discretization resolutions. Both models were trained on data with 100 time steps per loading cycle. The RNN model exhibits strong resolution dependence, with prediction error increasing as resolution deviates from the training resolution. In contrast, HANO maintains consistently low errors across all tested resolutions.
@@ 4.4 Generalization under multiple loading cycles
In the previous section, the training and validation sets consisted of two complete loading–unloading cycles. To assess the HANO model’s ability to extrapolate beyond its training range, we tested additional scenarios involving three, four, and five cycles.
@@
4.5 Robustness to noisy data
Real experimental measurements of strain and stress are inevitably affected by sensor noise. To verify whether the proposed HANO method can effectively filter out such noise and maintain accurate plasticity predictions, we conducted a experiment by superimposing random noise onto the stress signals and comparing the resulting predictions against the noise-free ground truth.
Following the elasto-plastic setup in Section 4.1, we added Gaussian noise to the stress at each increment:
@@ 4.5 Robustness to noisy data
where $\alpha=10\%$ is the noise ratio, and $\sigma_{\mathrm{std}}$ denotes the standard deviation of the underlying noise-free stress signals. After injecting this noise, we used the resulting noisy stress as part of the input to HANO, and quantified the prediction errors with respect to the noise-free reference. Figure 9 shows a representative noisy stress path (using $\alpha=10\%$), the HANO prediction, and the corresponding noise-free reference curve. As seen in the figure, the high-frequency components of the noise are strongly suppressed, and the predicted curve runs smoothly through the centre of the noisy envelope. This behaviour reflects the operator’s projection effect: the network maps the noisy input back onto the learned physical response manifold, acting as an implicit filter. Quantitatively, the NRMSE is around 0.015, demonstrating only a mild reduction in accuracy compared to the noiseless case.
@@
Overall, these findings suggest that HANO is inherently robust to sensor noise: it learns an operator mapping that respects the governing material laws, rather than overfitting to noisy observations. In practical applications where imperfect sensor data are unavoidable, this robustness ensures reliable stress predictions from potentially fluctuating signals.
@@ 5.2 Study of attention placement in HANO architectures
The first variant, HANO1, represents a naive introduction of self-attention, where the self-attention block is applied directly to the original input feature map. The intention here is to allow attention-driven refinements to guide the multi-scale feature extraction from the earliest stages of the network. The second variant, HANO2, follows a structure similar to that proposed in [51], which shows improved prediction accuracy over the vanilla FNO in the context of spatial physics simulations. In this setup, the self-attention block is applied to the feature map after the projection operator, and the resulting attention features are fused with outputs from both the Fourier layer and a linear transformation. Lastly, HANO3 introduces a self-attention module serially integrated with each U-Fourier layer, allowing both the downsampling and upsampling stages to benefit from attention-driven refinement of the feature maps. This third variant, HANO3, serves as the default architecture in this work, as illustrated in Figure 2 and employed in the modeling tasks presented in Section 4.
@@ 5.3 Comparison with Baseline Methods: FNO, UFNO, and HANO
Having established the importance of attention-placement strategies, we next compare the predictive performance of the proposed neural operator architecture against baseline models. Specifically, we contrast the standard Fourier Neural Operator (FNO), the U-Net-enhanced FNO (UFNO), and the top-performing HANO variant, HANO3 to illustrate how U-Net components and self-attention mechanisms improve stress predictions in history-dependent materials. For brevity, we hereafter refer to HANO3 simply as HANO.
@@
To validate that the observed improvements of HANO generalize beyond a single loading history, we evaluate three additional test paths randomly selected from the unseen test set Ge and Tagarielli [17], as shown in Figure 12. Across all cases, HANO consistently demonstrates high fidelity, accurately and robustly capturing both sharp transition features and the overall stress evolution.
These observations underscore the importance of architectural innovations beyond merely increasing model complexity, particularly for the challenging task of predicting long-term stress evolution in materials with strong history dependence. We also remark that while our focus in this Section is on the comparative analysis of different neural operator architectures, all neural operator models tested here outperform the standard neural networks originally used in Ge and Tagarielli [17].
@@ 6 Conclusion
This study demonstrates that the proposed HANO framework effectively addresses the two key dependency challenges commonly encountered in data-driven constitutive modeling. These findings underscore the importance of accounting for such factors to achieve robust generalization and faithful representation of irreversible material behavior. Looking forward, several extensions will be pursued to broaden HANO’s capabilities as a versatile framework for modeling complex inelastic materials. Owing to its end-to-end learning structure, the HANO-based constitutive model can be seamlessly integrated into differentiable numerical solvers to enable gradient-based inverse analysis, design optimization, and real-time simulation within a unified computational framework [64, 26, 12]. Moreover, HANO’s flexible architecture and input–output structure allow it to accommodate a wide range of inelastic material behaviors, including rate-dependent responses (e.g., viscoelastic and viscoplastic materials) and heterogeneous composite systems. While the present study focuses on purely data-driven modeling, HANO also lends itself to incorporating physics-based constraints and knowledge [44, 21, 27, 60], which may further enhance generalizability and predictive accuracy for complex inelastic materials, particularly under sparse or noisy data conditions.

### v2

@@ 4.2.1 Results and discussion
[…]
@@
By incorporating both strain and stress information and employing the designed neural operator framework, the proposed HANO approach effectively overcomes the zero-stress initialization constraints that limit conventional RNN-based models. This suggests its superior capacity for learning the hidden dynamics that govern path-dependent, irreversible behavior, enabling accurate stress predictions from any intermediate point along the loading history, even when initial data segments are missing. While this study primarily focuses on demonstrating HANO’s ability to address methodological limitations of existing approaches , we also conduct parameter-matched comparisons with RNN-based models to ensure fair evaluation, as detailed in D. Even under matched parameter counts, HANO demonstrates superior performance on both test scenarios, validating the effectiveness of the proposed architectural design.
@@ 4.3.2 Effect of discretization resolution in testset
[…]
Figure 4: Comparison of prediction errors (NRMSE) for RNN2 and HANO approaches with varying discretization resolutions in test dataset. Both models were trained on data with 100 time steps per loading cycle. The RNN model exhibits strong resolution dependence, with prediction error increasing as resolution deviates from the training resolution. In contrast, HANO maintains consistently low errors across all tested resolutions.
@@ 4.3.3 Effect of discretization resolution in training data
To assess the training and convergence behaviors of the proposed HANO framework, we provide the corresponding training and validation loss curves in E.
@@ 4.4 Generalization under multiple loading cycles
In this section, we assess the HANO model’s ability to extrapolate beyond its training range. As shown in Table 1, the baseline HANO model was trained on the original fixed-resolution dataset from Section 4.1 (100 time steps per cycle, 2 loading cycles). To evaluate its generalization capability, we tested additional scenarios involving three, four, and five cycles.
@@
4.5 Robustness to noisy initial condition
Real experimental measurements of strain and stress are inevitably affected by sensor noise. To verify whether the proposed HANO method can effectively filter out such noise and maintain accurate plasticity predictions, we conducted an experiment by superimposing random noise onto the initial stress signals and comparing the resulting predictions against the noise-free ground truth.
Following the elasto-plastic setup in Section 4.1, we added Gaussian noise to the $k$ initial stress inputs:
@@ 4.5 Robustness to noisy initial condition
where $\alpha=10\%$ is the noise ratio, and $\sigma_{\mathrm{std}}$ denotes the standard deviation of the underlying noise-free stress signals. After injecting this noise, we used the resulting noisy stress as part of the input to HANO, and quantified the prediction errors with respect to the noise-free reference. We note that due to HANO’s autoregressive prediction mechanism, noise is only added to the stress components within the initial history window. Once HANO begins making predictions, subsequent stresses are generated by the model itself rather than being provided as external data.
Figure 9 shows a representative stress-strain path with noisy input (using $\alpha=10\%$), the HANO prediction, and the corresponding noise-free reference curve. As seen in the figure, the high-frequency components of the noise are strongly suppressed, and the predicted curve runs smoothly through the centre of the noisy envelope. This behaviour reflects the operator’s projection effect: the network maps the noisy input back onto the learned physical response manifold, acting as an implicit filter. Quantitatively, the NRMSE is around 0.015, demonstrating only a mild reduction in accuracy compared to the noiseless case.
@@
Overall, these findings suggest that HANO is inherently robust to initial sensor noise: it learns an operator mapping that respects the governing material laws, rather than overfitting to noisy observations. In practical applications where imperfect sensor data are unavoidable, this robustness ensures reliable stress predictions from potentially fluctuating signals.
@@ 5.2 Study of attention placement in HANO architectures
The first variant, HANO1, represents a naive introduction of self-attention, where the self-attention block is applied directly to the original input feature map. The intention here is to allow attention-driven refinements to guide the multi-scale feature extraction from the earliest stages of the network. The second variant, HANO2, follows a structure similar to that proposed in [56], which shows improved prediction accuracy over the vanilla FNO in the context of spatial physics simulations. In this setup, the self-attention block is applied to the feature map after the projection operator, and the resulting attention features are fused with outputs from both the Fourier layer and a linear transformation. Lastly, HANO3 introduces a self-attention module serially integrated with each U-Fourier layer, allowing both the downsampling and upsampling stages to benefit from attention-driven refinement of the feature maps. This third variant, HANO3, serves as the default architecture in this work, as illustrated in Figure 2.
@@ 5.3 Comparison with Baseline Methods: FNO, UFNO, and HANO
Having established the importance of attention-placement strategies, we next compare the predictive performance of the proposed neural operator architecture against baseline models. Specifically, we contrast the standard Fourier Neural Operator (FNO), the U-Net-enhanced FNO (UFNO), and the top-performing HANO variant, HANO3 to illustrate how U-Net components and self-attention mechanisms improve stress predictions in history-dependent materials. For brevity, we hereafter refer to HANO3 simply as HANO. For comparison, the baseline models are configured as follows: FNO consists of 6 sequential Fourier layers with a channel width of $w=64$ and spectral resolution of 10 Fourier modes. UFNO comprises 3 standard Fourier layers followed by 3 U-Fourier layers (each incorporating parallel U-Net pathways), also with $w=64$ and 10 modes. Both baseline models use the same lifting and projection operators as HANO.
@@
To further evaluate the predictive performance of different neural operators with comparable model complexity, a parameter-matched comparison is conducted in Table 6 by enlarging the baseline models through increased channel widths: FNO-Large (width 72, $\sim$545k parameters) and UFNO-Large (width 68, $\sim$549k parameters). As shown in Table 6, despite having similar parameter counts, HANO (21.09% NRMSE) still outperforms both FNO-Large (33.76%) and UFNO-Large (24.12%). The enlarged FNO and UFNO models yield only marginal improvements, indicating that the performance gains of HANO primarily originate from the attention-enhanced U-Fourier architecture rather than from increased model size.
To validate that the observed improvements of HANO generalize beyond a single loading history, we evaluate three additional test paths randomly selected from the unseen test set [18], as shown in Figure 12. Across all cases, HANO consistently demonstrates high fidelity, accurately and robustly capturing both sharp transition features and the overall stress evolution.
These results underscore the importance of architectural innovations beyond simply increasing model complexity, especially for predicting long-term stress evolution in strongly history-dependent materials. The key contribution of this work is the development of an autoregressive neural operator framework for path-dependent materials, which eliminates the need for latent hidden states used in RNN-based models while delivering substantial performance gains. Although this is not an exhaustive architectural study, HANO could be further enhanced through additional refinements.
We also remark that while our focus in this Section is on the comparative analysis of different neural operator architectures, all neural operator models tested here outperform the standard neural networks originally used in Ge and Tagarielli [18].
@@
Model Parameter Count Performance (NRMSE)
FNO-Large 545,018 33.76%
UFNO-Large 548,632 24.12%
HANO 539,818 21.09%
Table 6: Parameter-matched comparison between FNO, UFNO, and HANO.
@@ 5.4 Sensitivity to history input length
An additional consideration worthy of discussion is the potential relationship between the minimum required input length and discretization resolution. For standard machine learning models such as RNNs, finer temporal resolution typically necessitates larger input windows to capture equivalent stress-strain variations, as these models rely on discrete representations where each time step captures less evolutionary information at higher resolutions. While HANO’s neural operator framework is designed to learn continuous dynamics independently of discretization (as demonstrated in Sections 4.3.2 and 4.3.3), the theoretical foundation established by Coleman and Gurtin [10] suggests that materials with memory fundamentally require a minimum observable window of external variables for accurate constitutive response prediction. This implies that even resolution-invariant models like HANO may encounter limitations when the physical time span covered by the history window becomes insufficient to capture the necessary historical information.
@@ 6 Conclusion
This study demonstrates that the proposed HANO framework effectively addresses the two key dependency challenges commonly encountered in data-driven constitutive modeling. These findings underscore the importance of accounting for such factors to achieve robust generalization and faithful representation of irreversible material behavior. Looking forward, several extensions will be pursued to broaden HANO’s capabilities as a versatile framework for modeling complex inelastic materials. Owing to its end-to-end learning structure, the HANO-based constitutive model offers potential to be integrated into differentiable numerical solvers to enable gradient-based inverse analysis, design optimization, and real-time simulation within a unified computational framework [71, 28, 12]. Moreover, HANO’s flexible architecture and input–output structure allow it to accommodate a wide range of inelastic material behaviors, including rate-dependent responses (e.g., viscoelastic and viscoplastic materials) and heterogeneous composite systems. While the present study focuses on purely data-driven modeling, HANO also lends itself to incorporating physics-based constraints and knowledge [48, 23, 29, 67], which may further enhance generalizability and predictive accuracy for complex inelastic materials, particularly under sparse or noisy data conditions.
Despite its demonstrated advantages, we acknowledge a potential limitation regarding the selection of the history window length $k$, a critical design parameter influenced by material complexity and loading patterns. Our analysis shows that HANO maintains robust performance across a range of $k$ values, with accuracy plateauing beyond a material-specific threshold. This behavior makes the framework practical for diverse applications where complete loading histories may not be available.
Acknowledgment
The authors gratefully acknowledge the constructive comments and suggestions provided by the anonymous reviewers, which have helped improve the quality and clarity of this manuscript. The work is partially supported by QH’s Startup Fund and Data Science Initiative (DSI) Seed Grant at the University of Minnesota.
Data availability
The training dataset used in this study will be made publicly available at: https://github.com/IntelligentMechanicsLab/HANO_inelasticity
@@ Appendix C Material parameters for dataset generation in anisotropic damage modeling
Appendix D Parameter-Matched Comparison with RNN Models
To ensure a fair comparison between HANO and RNN-based models, we conducted additional experiments with parameter-matched configurations. Following the experimental setup described in Section 4.2, we used the same 1D elastoplastic material model with identical training and testing datasets to evaluate model performance under controlled conditions.
[…]
All models were trained using identical protocols, including progressive scheduled sampling and noise injection strategies as described in Section 3.3.3. The evaluation was conducted on the same two test sets: Testset I (full loading history with undeformed start) and Testset II (partial loading history with deformed start). Table 8 presents the NRMSE results for all four models.
[…]
Parameters 77,121 77,185 455,685 77,089
Testset I 0.030 0.023 0.006 0.009
Testset II 0.359 0.106 0.004 0.008
Table 8: Parameter-matched comparison of the NRMSE error between RNN variants and HANO models for different test sets.
[…]
Appendix E Training and Validation Losses
The training and validation losses corresponding to the experiments in Section 4.3.3 are provided in this Appendix. Figures 14 shows the evolution of the Mean Squared Error (MSE) loss during training for the HANO model and the RNN2 baseline, respectively. The HANO model exhibits stable and efficient training behavior, with no signs of overfitting or underfitting. Both training and validation losses converge to similar final values around $10^{-4}$, indicating that the model has sufficient capacity to learn the underlying patterns while generalizing well to unseen data. In the initial training phase (approximately the first 200 epochs), there is a noticeable gap between training and validation losses, which is primarily attributed to the teacher forcing mechanism used in the progressive training strategy. As the scheduled sampling gradually transitions from teacher forcing to self-reliant prediction, the training and validation curves align more closely, demonstrating the effectiveness of this training approach.
The HANO model shows smooth and stable convergence throughout the training process. The training loss decreases consistently from approximately $10^{-1}$ to $10^{-4}$ over 500 epochs, while the validation loss follows a similar trajectory with minimal fluctuation. The model reaches stable convergence around epoch 400, with both losses plateauing at similar low values, confirming robust learning without overfitting.
In contrast, the RNN2 baseline exhibits different training dynamics. While the training loss converges smoothly to approximately $10^{-3}$, there is a persistent gap between training and validation losses throughout the training process. The validation loss shows more fluctuation compared to HANO, particularly in the later stages of training, suggesting less stable generalization behavior. The final validation loss remains approximately one order of magnitude higher than HANO’s final loss, indicating inferior performance.
Figure 14: Training and validation loss curves for the HANO and RNN2 model on the variable-resolution elastoplastic material dataset described in Section 4.3.3.